UNUSUAL BONDING MECHANISMS IN BERYLLIUM COMPOUNDS

IVAN O. ANTONOV, BEAU J. BARKER, VLADIMIR E. BONDYBEY, JEREMY M. MERRITT and \underline{MICHAEL C. HEAVEN}, Department of Chemistry, Emory University, Atlanta, GA 30322.

Metal clusters and small molecules that involve Be exhibit unusual bonding characteristics. For example, the Be_{2} dimer exhibits a weak bond, short equilibrium bond length and an atypical potential energy curve. Although the bond is weak, the dissociation energy of Be_{2} exceeds that of Mg_{2} by a factor of two. Be_{2}O is more strongly bound than would be expected on the basis of simple molecular orbital theory, and higher level theoretical calculations indicate that it has a highly multi-reference ground state. Indeed, calculations for many Be species indicate multi-reference character. To obtain results that are even qualitatively correct often requires the correlation of all electrons and the use of basis sets that include at least triple excitations. Consequently, spectroscopic data for these molecules provide a valuable set of benchmarks for the evaluation of highly correlated, multi-reference computational techniques. We are currently exploring the characteristics of Be compounds using electronic spectroscopy and photoionization techniques. The species being investigated include Be_{2}O, BeC and BeOH. Experimental results and theoretical calculations for these molecules will be presented, along with a progress report for studies of larger Be clusters.